Say that you purchase a house for $192,000 by getting a mortgage for $170,000 and paying a down payment of $22,000. If you get a 20-year mortgage with an interest rate of 6 percent, what are the monthly payments? (Round your final answer to 2 decimal places.)

Payment: ?

What would the loan balance be in five years? (Use a payment value rounded to 2 decimal places. Round your final answer to 2 decimal places.)

Loan Balance: ?

If the house appreciates at 3 percent per year, what will be the value of the house in five years? (Round your final answer to 2 decimal places.)

Future value: ?

How much of this value is your equity? (Use intermediate values rounded to 2 decimal places. Round your final answer to 2 decimal places.)

Equity: ?

Please answer all the parts correctly. THANK YOU!
Answer #1

1.
=170000*(6%/12)/(1-1/(1+6%/12)^(12*20))=1217.93

2.
=170000*(1+6%/12)^(12*5)-1217.93/(6%/12)*((1+6%/12)^(12*5)-1)=144329.51

3.
=192000*(1+3%)^5=222580.62

4.
=222580.62-144329.51=78251.11